What is the index of hydrogen deficiency of the following molecule? 0 6.5 0 0 AN 0 D
Answer #1

Dro Index of Mijdragen deficiency = f* (?e+2 – #=X+ N) C= No of C alems of Wys No. of H clams - 13 x = No. of Kalagens: 0 N